OpenStudy (anonymous):

The slope of a line is 0, and the y-intercept is 6. What is the equation of the line written in slope-intercept form? y = 6 y = 6x y = x + 6 c

OpenStudy (anonymous):

remember that slope intercept form means y = mx + b where m is the slope of the line and b is the y-intercept. Since you know m = 0 and b = 6 you can just plug these into y = mx + b and simplify.

OpenStudy (anonymous):

if my awnser is wrong i said C

OpenStudy (jdoe0001):

what did you get after you plugged in the values as @Twis7ed suggested?

OpenStudy (anonymous):

Yes it's wrong, do what I said and check what you get

OpenStudy (anonymous):

B

OpenStudy (anonymous):

What work did you do to get that answer??

OpenStudy (anonymous):

y=mx+b.....so y=0x+6 rite

OpenStudy (anonymous):

so y=6 ??

OpenStudy (jdoe0001):

ahemm sounds more like fishing for a letter

OpenStudy (anonymous):

so y=6 ??

OpenStudy (jdoe0001):

slope is rise/run rise is 0, how do you get 0? only if the 1st number and the 2nd number match so whatever - whatever = 0 if x = 3, y = 6 x = 5, y = 6 x = 1,000,000 y = 6  rise = 6-6 = 0 slope = 0/anything = 0 tis a horizontal line

OpenStudy (anonymous):

yeah so A

OpenStudy (anonymous):

??????????????????????

OpenStudy (jdoe0001):

heh yes, you're correct, the equation is y = 6